OpenStudy (anonymous):

Solving Equations with radicals 2(sqrt of x+3)=10

OpenStudy (zzr0ck3r):

\(2\sqrt{x+3}=10\) divide by 2 \(\sqrt{x+3}=5\) square both sides \(x+3=5^2=25\) so \(x = ?\)